OpenStudy (anonymous):

A sporting goods store has ordered x tents that cost $105 each. The shipping for the entire order is $160. The store can sell the tents for $215 each. find the formula for the profit function, p(x) in terms of x.

OpenStudy (anonymous):

(-105x -160) +215x = p(x)

OpenStudy (anonymous):

I'm tired. is that right? lol

OpenStudy (anonymous):

weird i has this P(x) = 215x - (105x + 160)

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

hmm what? just distribute your -1 in.. h a ha

OpenStudy (anonymous):

110x -160 = p(x)

OpenStudy (anonymous):

noo i was just be a clutz, has dislexia

OpenStudy (anonymous):

being*